OpenStudy (anonymous):

Evaluate the expression when a = 5 and b = 6 b +9a

OpenStudy (whpalmer4):

Hard to do when there's no expression provided :-)

OpenStudy (whpalmer4):

Okay, write "6" wherever you "b" and "5" wherever you see "a" in the expression. What does that equal?

OpenStudy (whpalmer4):

wherever you "b" -> wherever you see "b"

OpenStudy (texaschic101):

substitute 5 in for a and 6 in for b. b + 9a = 6 + 9(5) = your answer

OpenStudy (anonymous):

it wouldn't be 6 +45?

OpenStudy (whpalmer4):

It would be!

OpenStudy (texaschic101):

then what would it be ?

OpenStudy (anonymous):

51

OpenStudy (whpalmer4):

See, not so hard!

OpenStudy (anonymous):

but when they say evaluate do i just write 6 + 95

OpenStudy (whpalmer4):

Think of it like you are following a recipe... "I need 3 of these, and 4 of those, and 1 of these..." No, you want to collapse it down as far as you can, which would be to 51 in this case.

OpenStudy (anonymous):

oh ok.Thank you, I'm bad at math,lol

OpenStudy (whpalmer4):

It's just a matter of having it explained to you the right way (so that you understand it), and practice.

OpenStudy (texaschic101):

I wasnt saying that 6 + 45 was your answer....I just didnt finish and was going to let you finish

OpenStudy (whpalmer4):

As Henry Ford said, "whether you think you can do a thing, or you think you can't, you're right" — you *can* learn this stuff.